# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

### 06 Nov 2024 v1

@@ Appendix A The LTI condition is the only freedom in TI Hamiltonians
In the main text, we used the property that if two local operators supported on $n$ sites, satisfy the following condition
$\sum_{i}q_{n,i}=\sum_{i}g_{n,i}$ (47)
then they are related as (eq. 10 of main text)
$q_{n,i}=g_{n,i}+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}.$ (48)
Now, we will formally derive this property. Let $X$ be an $n$-body operator. Claim:
$\sum_{i}\tau_{i}(X)=0\iff\exists A\text{ such that }X=A\otimes\mathbb{I}-\mathbb{I}\otimes A.$ (49)
Here $\tau_{i}(X)$ denotes the translation of operator $X$ by $i$ sites.
Proof: For the direction $\mathrm{RHS}\Rightarrow\mathrm{LHS}$, notice that the terms $A\otimes\mathbb{I}-\mathbb{I}\otimes A$ cancel telescopically, meaning that
$\tau_{i}(A\otimes\mathbb{I}-\mathbb{I}\otimes A)=0.$ (50)
It remains to prove the other implication, i.e.,
$\tau_{i}(X)\Rightarrow X=A\otimes\mathbb{I}-\mathbb{I}\otimes A.$ (51)
To prove this, we take an orthonormal basis consisting of strings of single site operators $S_{i_{1},i_{2},\ldots,i_{n}}:=\sigma_{i_{1}}\otimes\sigma_{i_{2}}\otimes\ldots\sigma_{i_{n}}$ with $\sigma_{0}$ proportional to the identity $\mathbb{I}$. Expanding $X$ in this basis,
@@
In the above, first consider the basis elements with $i_{1}\neq 0$ and $i_{n}\neq 0$ and observe that in
$\textrm{Tr}\left(\sum_{j\in\mathbb{Z}}\tau_{j}(X)S_{i_{1},i_{2},\ldots,i_{n}}\right),$ (53)
we get that only one term in the sum survives, i.e. the term with $j=1$. Thus, $X$ satisfies
$X_{i_{1},i_{2},\ldots,i_{n}}=0\text{ for }i_{1},i_{n}\neq 0\text{ and for all }i_{2},\ldots,i_{n-1}=0,\ldots,d^{2}-1$ (54)
In the next step, we observe that $S_{0,i_{2},\ldots,i_{n}}$ with $i_{2}\neq 0$ and $i_{n}\neq 0$, which implies
$X_{0,i_{2},\ldots,i_{n}}+X_{i_{2},\ldots,i_{n},0}=0,$ (55)
i.e.
$X_{0,i_{2},\ldots,i_{n}}=-X_{i_{2},\ldots,i_{n},0},$ (56)
showing that $X$ has components $A^{(2)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(2)}$ with $A^{(2)}=\sum X_{i_{2},\ldots,i_{n},0}S_{i_{2},\ldots,i_{n}}$.
Continuing as before, we now observe that for $S_{0,0,i_{3},\ldots,i_{n}}$ with $i_{3}\neq 0$ and $i_{n}\neq 0$, we get
$X_{0,0,i_{3},\ldots,i_{n}}+X_{0,i_{3},\ldots,i_{n},0}+X_{i_{3},\ldots,i_{n},0,0}=0.$ (57)
Defining the $(n-2)$-body operators $B:=\sum_{i_{3},\ldots,i_{n}}X_{i_{3},\ldots,i_{n},0,0}S_{i_{3},\ldots,i_{n}}$, and $C:=\sum_{i_{3},\ldots,i_{n}}X_{0,i_{3},\ldots,i_{n},0}S_{i_{3},\ldots,i_{n}}$, we note that $X$ has contributions
$X=\ldots+$ $\sum_{i_{3},\ldots,i_{n}}X_{i_{3},\ldots,i_{n},0,0}S_{i_{3},\ldots,i_{n}}\otimes\mathbb{I}\otimes\mathbb{I}+$ (58)
$\sum_{i_{3},\ldots,i_{n}}X_{0,i_{3},\ldots,i_{n},0}\mathbb{I}\otimes S_{i_{3},\ldots,i_{n}}\otimes\mathbb{I}+$
$\sum_{i_{3},\ldots,i_{n}}(-X_{i_{3},\ldots,i_{n},0,0}-X_{0,i_{3},\ldots,i_{n},0})\mathbb{I}\otimes\mathbb{I}\otimes S_{i_{3},\ldots,i_{n}}=$
$=$ $\ldots+B\otimes\mathbb{I}\otimes\mathbb{I}+\mathbb{I}\otimes C\otimes\mathbb{I}-\mathbb{I}\otimes\mathbb{I}\otimes(B+C)$
Now defining $B^{\prime}=B$, $C^{\prime}=B+C$, we can rewrite the last line in eq. 58 as
$X=\ldots+B^{\prime}\otimes\mathbb{I}\otimes\mathbb{I}+\mathbb{I}\otimes(C^{\prime}-B^{\prime})\otimes\mathbb{I}-\mathbb{I}\otimes\mathbb{I}\otimes(C^{\prime})=\ldots+A^{(3)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(3)}$ (59)
with $A^{(3)}:=B^{\prime}\otimes\mathbb{I}+\mathbb{I}\otimes C^{\prime}$.
Repeating the same procedure for all $k\leq n$: Let $i_{k},i_{n}\neq 0$ and take the trace of $\sum_{j}\tau_{j}(X)$ with $S_{0,0,\ldots,0,i_{k},\ldots,i_{n}}$. Similarly as in the eq. 57 we obtain a condition which involves $k$ types of coefficients of $X$:
$X_{i_{k},\ldots,i_{n},0,\ldots,0}+X_{0,i_{k},\ldots,i_{n},\ldots,0}+X_{0,\ldots,i_{k},\ldots,i_{n},0}+\ldots+X_{0,\ldots,i_{k},\ldots,i_{n}}=0,$ (60)
which lead to contributions to $X$ of the form:
$X=\ldots+$ $T_{1}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$ (61)
$\mathbb{I}\otimes T_{2}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$
@@
$\mathbb{I}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots T_{k},$
where each $T_{i}$ is an $(n-k+1)$-body operator. Now for $l=1,\ldots,k-1$, we define the terms $T_{l}^{\prime}=\sum_{i=1}^{l}T_{i}$ and use $T_{k}=-\sum_{i=1}^{k-1}T_{i}=-T^{\prime}_{k-1}$ from eq. 60 to rewrite the terms in eq. 61 as $A^{(k)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(k)}$ where
$A^{(k)}:=[T_{1}^{\prime}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+\mathbb{I}\otimes T_{2}^{\prime}\otimes\ldots\mathbb{I}+\ldots\mathbb{I}\otimes\mathbb{I}\otimes\ldots T_{k-1}^{\prime}].$ (62)
Finally collecting all the coefficients of $X$, we obtain that $X=A\otimes\mathbb{I}-\mathbb{I}\otimes A$, where $A=\sum_{k=1}^{n}A^{(k)}$.
Appendix B First excited state is the optimal solution $\rho$ of the dual program
We take the dual objective function of the exact dual SDP eq. 15, and expand it in the eigenbasis of $H$,
[…]
$=\sum_{j}E_{j}^{2}\rho_{jj}$ (64)
$=\Delta^{2}\rho_{11}+\sum\limits_{j\geq 2}E_{j}^{2}\rho_{jj}.$ (65)
Similarly, expanding the constraint,
[…]
$=\Delta\rho_{11}+\sum\limits_{j\geq 2}E_{j}\rho_{jj}=1$ (67)
[…]
$=\Delta+\sum\limits_{j\geq 2}(E_{j}^{2}-\Delta E_{j})\rho_{jj},$ (69)
Since $E_{j}^{2}-\Delta E_{j}\geq 0\quad\forall\,j\geq 2$, the minimum is attained for $\rho_{jj}=0\quad\forall\,j\geq 2$. At the optimal $\rho^{*}$,
[…]
Now using positivity constraint on $\rho$, we find that
$\rho_{ij}=0\quad\forall\,\,i\neq j.$ (71)
which shows that the dual variable converges to the reduced marginals of the first excited state.
Appendix C Proof for inclusion of finite-size based methods
@@
In the following subsections, we derive $q_{n}(\delta)$ for the bound $\delta$ given by (1) Knabe bound [15], (2) Gosset–Mozgunov bound [20], (3) Martingale bound [14]. Following the assumptions of these methods, in the proof, we focus on the case of nearest neighbor Hamiltonians, and assume that the local terms are projectors, i.e. $h_{i}^{2}=h_{i}$.
C.1 Knabe bound eq. 22
@@
$h_{n,i}=\sum_{j=i}^{i+n-2}h_{j,j+1}.$ (74)
where the index $i$ denotes the index of the starting site. Note that this Hamiltonian is supported on sites $i$ to $i+n-1$. The open boundary Hamiltonian satisfies the following relation,
@@
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right).$ (76)
@@
$Q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $=H+\sum_{i}\left(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}\right)-\Delta_{\mathrm{Knabe}}(n)H,$ (81)
@@
$=\left(\frac{1-\epsilon_{n}}{n-2}\right)(n-1)H+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (83)
Now, we notice that $(n-1)H$ can be expanded as
@@
which when summed over $i$ on both sides, recover the same global operator. For example, in the first replacement, the two sides are locally related by an operator $Y$ supported on $n-1$ sites, which satisfies the equation,

### 31 Mar 2026 v2

@@ Appendix A The LTI condition is the only freedom in TI Hamiltonians
In the main text, we used the property that if $g_{n}$ and $q_{n}$ generate the same translation invariant operator on a spin chain with $m$ sites and periodic boundary conditions, i.e., if
$\sum_{i=1}^{m}\tau_{i}(q_{n})=\sum_{i=1}^{m}\tau_{i}(g_{n}),$ (49)
where $\tau_{i}$ is the translation operator such that $\tau_{i}(g_{n})$ acts on sites $\{i,\ldots,i+n-1\}$, then they are related as (eq. 10 of main text):
$q_{n}=g_{n}+A\otimes\mathbb{I}-\mathbb{I}\otimes A,$ (50)
for some (n1)(n-1)-body operator $A$. We now prove this claim: Let $X$ be an $n$-body operator and let $m\geq 2n-1$, then $\sum_{i=1}^{m}\tau_{i}(X)=0$ holds iff there exists an (n1)(n-1)-body operator $A$ such that $X=A\otimes\mathbb{I}-\mathbb{I}\otimes A$.
Proof: For the implication $X=A\otimes\mathbb{I}-\mathbb{I}\otimes A$ $\Rightarrow$ $\sum_{i=1}^{m}\tau_{i}(X)=0$, notice that the terms $\tau_{i+1}(A\otimes\mathbb{I})$ and $\tau_{i}(\mathbb{I}\otimes A)$ appear with opposite signs and thus the whole sum cancels telescopically, meaning that
$\sum_{i}\tau_{i}(A\otimes\mathbb{I}-\mathbb{I}\otimes A)=0.$ (51)
It remains to prove the other implication. To prove this, we take an orthonormal basis consisting of strings of single site operators $S_{i_{1},i_{2},\ldots,i_{n}}:=\sigma_{i_{1}}\otimes\sigma_{i_{2}}\otimes\ldots\sigma_{i_{n}}$ with $\sigma_{0}$ proportional to the identity $\mathbb{I}$. We write out the coefficients of $X$ in this basis,
@@
First consider the basis elements with $i_{1}\neq 0$, $i_{n}\neq 0$, and $i_{2},\ldots,i_{n-1}$ arbitrary, and compute the trace of $S_{i_{1},i_{2},\ldots,i_{n}}\otimes\mathbb{I}_{n+1,\ldots,m}$ with $\sum_{j}\tau_{j}(X)$. We see that only one term in the sum survives (the one with $j=1$):
$0=\textrm{Tr}\left(\sum^{m}_{j=1}\tau_{j}(X)S_{i_{1},i_{2},\ldots,i_{n}}\right)=X_{i_{1},i_{2},\ldots,i_{n}},\;\;\text{for }i_{1}\neq 0,i_{n}\neq 0,\text{and }i_{2},\ldots,i_{n-1}=0,\ldots,d^{2}-1.$ (53)
We thus see that the part of $X$ which has non trivial support on sites 1 and $n$ must vanish.
Now let $n\geq k\geq 2$ and let $i_{j}=0$ for all $j<k$ and $i_{k},i_{n}\neq 0$, and let $i_{k+1},\ldots,i_{n-1}$ be arbitrary. Taking the trace of $\sum_{j}\tau_{j}(X)$ with $S_{0,0,\ldots,0,i_{k},\ldots,i_{n}}\otimes\mathbb{I}_{n+1,\ldots,m}$ we get that
$X_{i_{k},\ldots,i_{n},0,\ldots,0}+X_{0,i_{k},\ldots,i_{n},\ldots,0}+\ldots+X_{0,\ldots,i_{k},\ldots,i_{n}}=0.$ (54)
We will now use this relation between the coefficients of $X$ corresponding to basis elements with a non trivial (nk+1)(n-k+1)-body operator somewhere in the string (eq. 54) to show that the part of $X$ with non trivial support on nk+1n-k+1 sites can be written as $A^{(n-k+1)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(n-k+1)}$ for some $A^{(n-k+1)}$. Let $X^{(n-k+1)}$, “the (nk+1)(n-k+1)-body part of XX”, be defined as
$X^{(n-k+1)}:=\sum_{\begin{subarray}{c}i_{k}\neq 0,i_{n}\neq 0,\\ i_{k+1},\ldots,i_{n-1}\end{subarray}}X_{i_{k},\ldots,i_{n},0,\ldots,0}S_{i_{k},\ldots,i_{n},0,\ldots,0}+X_{0,i_{k},\ldots,i_{n},\ldots,0}S_{0,i_{k},\ldots,i_{n},\ldots,0}+\ldots+X_{0,\ldots,i_{k},\ldots,i_{n}}S_{0,\ldots,i_{k},\ldots,i_{n}}.$ (55)
For $l=1,\ldots,k$ define
$T^{(n-k+1)}_{l}:=\sum_{\begin{subarray}{c}i_{k}\neq 0,i_{n}\neq 0,\\ i_{k+1},\ldots,i_{n-1}\end{subarray}}X_{\underset{l-1\text{ times}}{0,\ldots\ldots,0},i_{k},\ldots,i_{n},\underset{k-l\text{ times}}{0,\ldots\ldots,0}}S_{i_{k},\ldots,i_{n}}\times d^{(k-1)/2},$ (56)
such that
$X^{(n-k+1)}=$ $T^{(n-k+1)}_{1}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$ (57)
$\mathbb{I}\otimes T^{(n-k+1)}_{2}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$
@@
$\mathbb{I}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots T^{(n-k+1)}_{k}$.
Now for $j=1,\ldots,k-1$, we define $T^{\prime(n-k+1)}_{j}=\sum_{l=1}^{j}T^{(n-k+1)}_{l}$, and
$A^{(n-k+1)}:=$ $T^{\prime(n-k+1)}_{1}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$ (58)
$\mathbb{I}\otimes T^{\prime(n-k+1)}_{2}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$
$\vdots$
$\mathbb{I}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots T^{\prime(n-k+1)}_{k-1}$.
Equation 54 implies $T^{(n-k+1)}_{k}=-\sum_{l=1}^{k-1}T^{(n-k+1)}_{l}=-T^{\prime(n-k+1)}_{k-1}$ from which we obtain that
$X^{(n-k+1)}=A^{(n-k+1)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(n-k+1)}.$ (59)
Since $X=\sum_{k=2}^{n}X^{(n-k+1)}$ (we have shown in the beginning that $X^{(n)}=0$), we have that $X=A\otimes\mathbb{I}-\mathbb{I}\otimes A$, with $A=\sum_{k=2}^{n}A^{(n-k+1)}$ as claimed.
Appendix B Proof for inclusion of finite-size based methods
@@
In the following subsections we explicitly derive $q_{n}(\delta)$ for the bound $\delta$ given by the Knabe bound [32] (section B.1) and the Martingale bound [44] (section B.3). For the Gosset–Mozgunov bound [24] we prove that such a generating term exists without explicitly computing it (section B.2).
Following the assumptions of these methods, in the proof, we focus on the case of nearest neighbor Hamiltonians, and assume that the local terms are projectors, i.e. $h_{i}^{2}=h_{i}$.
B.1 Knabe bound eq. 23
@@
$h_{n,i}=\sum_{j=i}^{i+n-2}h_{j}.$ (62)
where the index $i$ denotes the index of the starting site. Note that this Hamiltonian is supported on sites $i$ to i+n1i+n-1. The open boundary Hamiltonian satisfies the following relation,
@@
$\Delta_{m}\geq\delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right).$ (64)
@@
$Q_{n}(\delta_{\mathrm{Knabe}}(n))$ $=H^{2}-\delta_{\mathrm{Knabe}}(n)H$ (69)
$=H+\sum_{i}\left(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}\right)-\delta_{\mathrm{Knabe}}(n)H,$ (70)
@@
$=\left(\frac{1-\epsilon_{n}}{n-2}\right)(n-1)H+\sum_{i}\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (72)
Now, we notice that (n1)H(n-1)H can be expanded as
@@
which when summed over $i$ on both sides, recover the same global operator. For example, in the first replacement, the two sides are locally related by an operator $Y$ supported on n1n-1 sites, which satisfies the equation,